OpenStudy (anonymous):

Find the equation of a line passing through (-2,5) and (3,4) in Ax+By=C

OpenStudy (shamil98):

First, find the slope, \[\large \frac{ y_2-y_1 }{ x_2-x_1 }\] \[\large \frac{ 4-5 }{ 3 + 2 } = -1/5\] Input your x and y coordinate from one of those to get b. 5 = (-1/5)(-2) + b 5 = 2/5 + b 4 3/5 = b your resulting equation is. \[\large y = \frac{ 1 }{ 5 }x +4\frac{ 3 }{ 5 }\] from there you need to get whole numbers and get x on the other side.

OpenStudy (shamil98):

\[\large -\frac{ 1 }{ 5 }x + y = \frac{ 23 }{ 5 }\] Now you need to multiply the equation by a number so that we get whole numbers. Let's multiply the equation by 5. \[\large 5 \times (\frac{ -1 }{ 5 }x + y = \frac{ 23 }{ 5 })\] you get \[\large \frac{ -5 }{ 5 }x + 5y = \frac{ 115 }{ 5 }\] reduce the fractions. \[\large -x + 5y = 23\] This is your standard form for your equation.